Shaypop:

M ‹ A = 100-x m < B = 80+x m < C = 110-3x m < D = 75+2x Quadrilateral ABCD is a parallelogram if both pairs of opposite angles are equal. Prove that quadrilateral ABCD is a parallelogram by finding the value of x.

Shaypop:

AZ:

so opposite angles are equal that means angle A = angle C angle B = angle D

AZ:

angle A is 100 - x angle C is 110 - 3x so set them equal to each other and solve for x 100 - x = 110 - 3x

Shaypop:

5

AZ:

Good! And similarly angle B = angle D 80 + x = 75 + 2x

Shaypop:

Same thing

AZ:

Exactly! So we found the value of x to prove that the quadrilateral is a parallelogram